what would be the equation for the following?

(a) solid carbon plus oxygen gas burns to produce gaseous carbon monoxide


(b) dissolved silver nitrate plus aqueous sodium chloride produces solid silver chlorid and aqueous sodium nitrate.
1
Expert's answer
2012-06-22T08:39:11-0400
a) 2 C (s) + O2 (g) = 2 CO (g)
b) AgNO3 (aq) + NaCl (aq) = AgCl (s) + NaNO3
(aq)
Ag(+) + NO3(-) + Na(+) + Cl(-) = AgCl (s) + Na(+) + NO3(-)

Ag(+) + Cl(-) = AgCl (s)